I need help on a word problem.
A 20 meter ladder and a 15 meter ladder were leaned against a building. The bottom of the longer ladder was 7 meters farther from the building than the bottomof the shorter ladder, but both ladders reached the same distance up the building. Find the distance the ladders reached up the building.
I don't even remember how to begin to do this.

Answer by stanbon(57940)

I need help on a word problem.
A 20 meter ladder and a 15 meter ladder were leaned against a building. The bottom of the longer ladder was 7 meters farther from the building than the bottomof the shorter ladder, but both ladders reached the same distance up the building. Find the distance the ladders reached up the building.
----------------
Draw the two right triangle.
longer hypotenuse = 20 meters; shorter hypotenuse = 15 meters
----
longer base = x+7 ; shorter base = x
----
height for both is the same
==================================
20^2 = h^2 + (x+7)^2
15^2 = h^2 + x^2
-----------------------
Subtract to solve for "x":
175 = (x^2+14x+49) - x^2
14x+48 = 175
14x = 127
x = 9.07 meters
x+7 = 16.07 meters
=========================
Cheers,
Stan H.
=============